methane is a fuel. it completely burns to form carbondioxide and water. when one mole of methane is burned 890 kj of energy is released
i don't know they've given these some symbols nd i don't how to solve them let me try writing them
H-C-H apart from that one H is over the C and the other one is below C than +2o=o gives you o=c=o+2 o and H on both the upper sides
1
Expert's answer
2012-10-11T09:29:04-0400
& H
I
H--- C----H + 2 O=O& ------> O=C=O + 2 H--O--H + Q (Q= 890kj)
I
& H